OpenStudy (anonymous):

What is the simplified form of (5x+4/x+7)-(2x+1/x+7)

OpenStudy (maheshmeghwal9):

\[(5x+4-2x-1)/(x+7)=(3x+3)/(x+7)=3(x+1)/(x+7)\]

OpenStudy (maheshmeghwal9):

First step for solving:) Since {-} should be multiplied with each term inside the bracket.